Red gold is a gold‑copper alloy used to make jewelry. A piece of jewelry made of red gold weighs 8.62 g

and has a volume of 0.475 cm

3

. Gold has a density of 19.3 g/cm

3

and copper has a density of 8.96 g/cm

3

. Calculate the percentage by mass of each metal in the jewelry. Assume the total volume of the jewelry is the sum of the volumes of the two metals it contains.


1
Expert's answer
2021-02-09T03:35:20-0500

19.3x + 8.96y= 8.62

x + y= 0.475.

x= 0,421947674

y= 0,0530523256.

19.3 x 0,421947674 x 100%/8.62 = 94,4732031 % of gold.

And 100-94,4732031 = 5,5267969%. Of Copper